Calculate the rms speed of helium atoms near the surface of the Sun at a temperature of about 6400 K. Express your answer to two significant figures and include the appropriate units. vrms v r m s = nothingnothing SubmitRequest Answer
Answer #1

v^2 = 3 KT /m

v^2 = 3* 1.38*10^-23* 6400/ ( 4*1.67*10^-27)

v = 6298 m/s

==========

do comment in case any doubt, will reply for sure..Goodluck